# Resolution of Simpson’s paradox via the common cause principle

### v1

@@ I Introduction
Figure 1: Directed acyclic graphs between random variables $A=(A_{1},A_{2})$, $B$ and $C$ involved in discussing Simpson’s paradox. The first and second graphs were studied in Refs. pearl ; pearl2 . They are discussed in section I and Appendix A. The third or fourth graphs are basic assumptions of this work; see (4). In the first graph, $B$ influences $A_{1}$ and $A_{2}$, but $B$ is not the common cause in the strict sense, because there is an influence from $A_{2}$ to $A_{1}$. A similar interpretation applies to the second graph. We emphasize that the joint probability $p(A_{1},A_{2},B)$ for the first and second graphs has the same form, i.e. such graphs are extra constructions involved for an additional interpretation of data. In contrast, the third and fourth graph imply a definite (but the same for both graphs) limitation on the joint probability $p(A_{1},A_{2},B,C)$, which is expressed by (4).
II Common cause principle and reformulation of Simpson’s paradox
The common cause for $A=(A_{1},A_{2})$ and $B$ means reich ; suppes
@@ II Common cause principle and reformulation of Simpson’s paradox
where this screening relation holds for all values assumed by $A_{1}$, $A_{2}$, $B$ and $C$. The times of these variables naturally hold $t_{C}<{\rm min}[t_{A_{1}},t_{A_{2}},t_{B}]$. The same (4) applies if $C$ causes $A$ and screens $A$ from $B$. Now $t_{B}<t_{C}<{\rm min}[t_{A_{1}},t_{A_{2}}]$. These two scenarios are shown in Fig. 1 as (resp.) the third and fourth graphs. Appendix A provides several examples for a causing (or screening) variable $C$ in the context of Simpson’s paradox.
Note from (4) that $C$ gets rid of the conditional dependence on $B$ in $p(A_{1},A_{2}|B,C)$. Thus, the correct way of looking at the association between $a_{1}$ and $a_{2}$ is to check the sign of $p(a_{1}|a_{2},C)-p(a_{1}|\bar{a}_{2},C)$ for all values of $C$. To support this point, we note that (4) has an important implication in the context of (1). (This implication generalizes the argument given in suppes .) Assume that $p(a_{2},b,c)>0$ for all values $c$ of $C$. Note from (4) that there exists an event $c$ such that $p(a_{1}|a_{2},b)\leq p(a_{1}|a_{2},b,c)=p(a_{1}|a_{2},c)$, and an event $c^{\prime}$ such that $p(a_{1}|a_{2},b)\geq p(a_{1}|a_{2},b,c^{\prime})=p(a_{1}|a_{2},c^{\prime})$. Hence, if conditioning over $b$ facilitates (hinders) the association between $a_{1}$ and $a_{2}$, then conditioning over $c$ ($c^{\prime}$) is not worse in this facilitation (hindering)  [^4] To deduce the first relation assume that $p(a_{1}|a_{2},b)<p(a_{1}|a_{2},b,c)=p(a_{1}|a_{2},c)$ for all $c$, multiply both parts by $p(a_{2},b,c)>0$, sum over $c$ and get contradiction $p(a_{1},a_{2},b)<p(a_{1},a_{2},b)$. Likewise for the second relation..
@@
Theorem: If $A_{1}$, $A_{2}$, $B$ and $C=\{c,\bar{c}\}$ are binary, and provided that (1) and (2, 3) are valid, all causes $C$ hold
@@
i.e. all $C$ holding (4) predict the same sign of association between $a_{1}$ and $a_{2}$ as (2, 3). The main idea of proving (5) is inverting (4), i.e. presenting it as
$p(a_{1}|a_{2},c)$ $=$ $\frac{p(\bar{b}|\bar{c})p(a_{1}|a_{2},b)p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(a_{1}|a_{2},\bar{b})p(\bar{b}|a_{2})}{p(\bar{b}|\bar{c})p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(\bar{b}|a_{2})}$ (6)
@@
$p(c|a_{2})$ $=$ $\frac{p(\bar{b}|\bar{c})p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(\bar{b}|a_{2})}{p(b|c)+p(\bar{b}|\bar{c})-1}=\frac{p(\bar{b}|\bar{c})+p(b|a_{2})-1}{p(b|c)+p(\bar{b}|\bar{c})-1}$ (8)
where unknown quantities $p(a_{1}|a_{2},c)$ and $p(c|a_{2})$ are represented via known ones (i.e. $p(A_{1},A_{2},B)$) and free parameters $p(B|C)$. Eqs. (7, 8) hold upon changing $a_{2}$ by $\bar{a}_{2}$. Eqs. (7, 8) are deduced in Appendix B via specific notations that should be useful when dealing with (4) more generally, i.e. for non-binary $C$.
The rest of the proof is algebraic but non-trivial. It also works out and employs constraints (13, 18) on Simpson’s paradox itself. Expanding both sides of (1),
@@
and using there (2, 3) we subtract both sides of (1) from each other and find:
$p(a_{1}|a_{2},b)+p(\bar{b}|a_{2})[p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b)]<p(a_{1}|\bar{a}_{2},b)+p(\bar{b}|\bar{a}_{2})[p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)].$ (11)
@@
Eqs. (9, 10) imply that for the validity of (13, 12) it is necessary to have $p(a_{1}|\bar{a}_{2},\bar{b})>p(a_{1}|a_{2},b)$, which together with (2, 3, 12) can be written as
$p(a_{1}|a_{2},\bar{b})>p(a_{1}|\bar{a}_{2},\bar{b})>p(a_{1}|a_{2},b)>p(a_{1}|\bar{a}_{2},{b}).$ (13)
Now (1, 9, 10) can be written as
$p(a_{1}|a_{2},b)+[1-p(b|a_{2})](p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))<p(a_{1}|\bar{a}_{2},b)+[1-p(b|\bar{a}_{2})](p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$ (14)
$p(a_{1}|a_{2},\bar{b})-p(b|a_{2})(p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))<p(a_{1}|\bar{a}_{2},\bar{b})-p(b|\bar{a}_{2})(p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$ (15)
where (14) and (15) are equivalent. Now (14) and (15) imply from (13) (resp.)
$[1-p(b|a_{2})](p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))<[1-p(b|\bar{a}_{2})](p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$ (16)
$p(b|a_{2})(p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))>p(b|\bar{a}_{2})(p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)).$ (17)
As checked directly, Eqs. (16, 17) lead to
@@
Now we return to (8) and assume there ${p(b|c)+p(\bar{b}|\bar{c})-1}<0$, which leads to ${p(\bar{b}|\bar{c})+p(b|a_{2})-1}<0$ from (8). Writing down from (8) the formula for $p(c|\bar{a}_{2})$ and making the same assumption we get ${p(\bar{b}|\bar{c})+p(b|\bar{a}_{2})-1}<0$. Now look at (7) and its analog obtained via $a_{2}\to\bar{a}_{2}$, and use there these two results together with (17, 18) and (13) to deduce the first inequality in (5) under assumption ${p(b|c)+p(\bar{b}|\bar{c})-1}<0$. It should be obvious that the second inequality in (5) holds under the same assumption since we nowhere used any specific feature of $c$ compared to $\bar{c}$.
@@
The rest is similar to the above: we proceed via (16, 18) and (13) and deduce (5) from (8), (19) and the analog of (19) obtained via $a_{1}\to\bar{a}_{2}$.
III Summary
We addressed Simpson’s paradox: the problem of setting up an association between two events $a_{1}$, $a_{2}$ given the lurking variable $B$. The paradox provides two plausible but opposite conclusions for the same situation; see (1) and (2, 3). Thus, there are three possible outcomes: either the first option is correct, the second option is correct, or none of them are correct.
We focus on cases, when there is a common cause $C$ for $B$ and $A=(A_{1},A_{2})$ (which combines $a_{1}$, $a_{2}$ and their complements). Alternatively, $C$ screens out $A$ from $B$. These cases include those in which there is no causal influence from $A$ to $B$, as well as from $B$ to $A$. Hence, correlations between $A$ and $B$ are to be explained via the common cause $C$, which is a statement of the common cause principle reich ; suppes . Now the association between $a_{1}$ and $a_{2}$ is to be decided by looking at $p(a_{1}|a_{2},c)$ for various values of $C$. This task is normally difficult given the fact that $C$ is frequently not fully known and is not observed. However, provided that $A_{1}$, $A_{2}$, $B$ and $C$ are binary, $p(a_{1}|a_{2},c)$ shows the same association as the option (2, 3) of Simpson’s paradox. In this sense, Simpson’s paradox is resolved in the binary situation, provided that the situation allows a binary cause or a binary screening variable.
@@ III Summary
Starting from a tertiary $C$, all three options of Simpson’s paradox become possible: there are common causes $C$ that support (1), those which support (2, 3), and eventually (random) cause variables $C={c_{1},c_{2},c_{3}}$ for which $p(a_{1}|a_{2},c_{i})-p(a_{1}|\bar{a}_{2},c_{i})$ has different signs for different values of $i=1,2,3$. Our numerical examples showing these possibilities are contained in Supplementnumerics . Hence, already for the tertiary cause one needs prior information on the common cause to decide on the solution of Simpson’s paradox. Alternatively, we can infer this unknown cause via e.g. one of methods proposed recently koc ; hovh2023 . It is not excluded that such inference methods will provide further information on the solution of Simpson’s paradox. We hope to discuss this problem elsewhere.
Acknowledgements.
We were supported by HESC of Armenia, grants 24FP-1F030 and 21AG-1C038.
Appendix A Examples of Simpson’s paradox
We collected several examples of the paradox that are scattered in the literature. We discuss them employing our notations in (13) emphasizing (whenever relevant) the existence of the common cause (or screening) variable $C$.
Example 1. Medical treatment lindley . $A_{1}=\{a_{1},\bar{a}_{1}\}$ (the target variable) is the recovery rate of medical patients: $a_{1}={\rm recovery}$, $\bar{a}_{1}={\rm no~{}recovery}$. $A_{2}=\{a_{2},\bar{a}_{2}\}$ refers to a specific medical treatment: $a_{2}={\rm treatment}$, $\bar{a}_{2}={\rm no~{}treatment}$. $B=\{b,\bar{b}\}$ is the sex of patients: $b={\rm male}$, $\bar{b}={\rm female}$. The times to which the random variables $A_{1}$, $A_{2}$ and $B$ refer clearly hold $t_{B}<t_{A_{2}}<t_{A_{1}}$. Ref. lindley argued for this example that a better method to decide on the association between $A_{1}$ and $A_{2}$ is (2, 3).
Given that this example holds Simpson’s paradox and assuming that it can be described via a time-ordered direct acyclic graph (TODAG), there is only one TODAG that accounts for it: $B\to A_{2}\to A_{1}\leftarrow B$; see the first graph in Fig. 1. Then $B$ causes both $A_{1}$ and $A_{2}$, and—according to lindley ; pearl ; pearl2 —it is natural to condition over it leading to the decision rule (2, 3).
It is possible as well that this situation is not described by any TODAG, because there is no direct causal influence between $A=(A_{1},A_{2})$ and $B$. Then one possible common cause for $A$ and $B$ is the genetic makeup.
Example 2. Plant yield lindley . $A_{1}=\{a_{1},\bar{a}_{1}\}$ (the target variable) is the yield of a single plant: $a_{1}={\rm high}$, $\bar{a}_{1}={\rm low}$. $A_{2}=\{a_{2},\bar{a}_{2}\}$ refers to the variety (color) of the plant: $a_{2}={\rm dark}$, $\bar{a}_{2}={\rm light}$. $B=\{b,\bar{b}\}$ refers to the height of the plant: $b_{1}={\rm tall}$, $\bar{b}={\rm low}$. The times hold $t_{A_{2}}<t_{B}<t_{A_{1}}$. Given that this situation holds Simpson’s paradox and assuming that it can be described via a time-ordered direct acyclic graph (TODAG), there is only one TODAG that accounts for it: $A_{2}\to B\to A_{1}\leftarrow A_{2}$; see the second graph in Fig. 1. Now since $B$ blocks one possible path from $A_{2}$ to $A_{1}$, Ref. pearl2 proposes not to condition over it suggesting that the correct decision rule is (1). This also coincides with the suggestion made in Ref. lindley concerning this example.
Again, the data itself does not assume any direct causal influence between $A=(A_{1},A_{2})$ and $B$, i.e. this example need not be described by any TODAG. One possible common cause $C$ for $A$ and $B$ is the genetic makeup.
Example 3. This example is taken from Ref. appleton . Here $A_{1}$ amounts to the survival of a person (i.e. reaching a certain age), the binary $A_{2}$ refers to smoker versus non-smoker, while $B$ means the age group. It turns out that the aggregated data (1) shows that smoking is beneficial for survival. This conclusion is reversed, once the age group $B$ is introduced, i.e. (2, 3) is employed instead of (1). The peculiar point of this example is that according to Ref. appleton , $B$ (i.e., the number of people of a certain age) was influenced by $A_{2}$. Hence this example is of the same type as Example 2 and could be described by the TODAG $A_{2}\to B\to A_{1}\leftarrow A_{2}$. However, marginalizing over $B$ (i.e. employing (1)) will lead to a wrong interpretation that smoking is good for survival.
Example 4. Snow tires provide cars with better traction in snowy and icy road conditions. However, nationally in the US, cars fitted with snow tires are more likely to have accidents in snowy and icy conditions norton . $A_{1}=\{a_{1}={\rm accident},\bar{a}_{1}={\rm no-accident}\}$, $A_{2}=\{a_{2}={\rm changed~{}tires},\bar{a}_{2}={\rm not~{}changed}\}$, $B=\{{\rm states}\}$. Here the choice of the state (warm or cold) has a direct causal link to accidents in winter conditions. Now snow tires tend to be fitted to cars only in snowy winter months and in states with colder weather. Cars in warmer months and in states with warmer weather are much less likely to have accidents in snowy and icy conditions. Plausibly, there is a random variable, $C=\{{\rm good~{}weather~{}conditions},{\rm bad~{}weather~{}conditions}\}$, which causes $A$, and screens $A$ from $B$: $p(A|CB)=p(A|C)$. The times are distributed as $t_{B}<t_{C}<t_{A_{2}}<t_{A_{1}}$.
Example 5. We need to choose between two hospitals 1 and 2: $A_{1}=\{a_{1}={\rm recovered},\bar{a}_{1}={\rm not~{}recovered}\}$, $A_{2}=\{a_{2}={\rm hospital~{}1},\bar{a}_{2}={\rm hospital~{}2}\}$, $B=\{{\rm first~{}half-year},{\rm second~{}half-year}\}$, $C=\{{\rm types~{}of~{}illness}\}$. Here we do not expect direct causal influence from $B$ to $A$, if (as we assume) the hospitals do not treat seasonal illnesses. We expect that $C$ causes $A$, and screens it from $B$.
Note that the data from which the probabilities for Simpson’s paradox are calculated is the number of patients $N(A_{1},A_{2},B)$ that came to the hospital. Simpson’s paradox does not exist—moreover, the decisions based on conditioning and on marginalization of $B$ are equivalent—if within each season the hospitals accept an equal number of patients: $\sum_{A_{1}}N(A_{1},A_{2},B)=\sum_{A_{1}}N(A_{1},A_{2},B)$ for each value of $B$. This creates a conceptual possibility for judging between the hospitals. This is however not realistic, because imposing on these hospitals an equal number of patients can disturb their usual (normal) functioning.
Example 6. Simpson’s paradox is realized when comparing scores of professional athletes, e.g. the batting averages of baseball players ross . Here $A_{1}$ refers to a score of an athlete in a game, e.g. $A_{1}={\rm high~{}score,low~{}score}$, $A_{2}$ denotes concrete athletes, while $B$ is the time-period (e.g. playing season). The causing variable $C$ can refer to the psychological and physical state of an athlete that influences his/her game success, and the number of games he/she participated in each season.

### v2

@@ 3 Common cause principle and reformulation of Simpson’s paradox
Figure 1: Directed acyclic graphs between random variables $A=(A_{1},A_{2})$, $B$ and $C$ involved in discussing Simpson’s paradox. The first and second graphs were studied in Refs. [13, 14]; see (8, 9). The third or fourth graphs are basic assumptions of this work; see (15). In the first graph, $B$ influences $A_{1}$ and $A_{2}$, but $B$ is not the common cause in the strict sense, because there is an influence from $A_{2}$ to $A_{1}$. A similar interpretation applies to the second graph. We emphasize that the joint probability $p(A_{1},A_{2},B)$ for the first and second graphs has the same form, i.e. such graphs are extra constructions employed for interpretation of data. In contrast, the third and fourth graph imply a definite (but the same for both graphs) limitation on the joint probability $p(A_{1},A_{2},B,C)$, which is expressed by (15).
3.1 Common cause and screening
The common cause for $A=(A_{1},A_{2})$ and $B$ means that there exists a random variable $C=\{c\}$ [34, 35]
@@ 3.1 Common cause and screening
$p(A_{1},A_{2},B)={\sum}_{c}\,p(A_{1},A_{2}|c)p(B|c),$ (16)
where (15) holds for all values assumed by $A_{1}$, $A_{2}$, $B$ and $C$, and where (16) follows from (15)  [^3] There are formulations of the common cause principle that look for (15) holding for certain events only and not for random variables [34, 35]. We do not focus on them. . The same (15) applies if $C$ causes $A$ and screens $A$ from $B$. These two scenarios are shown in Fig. 1 as (resp.) the third and fourth graphs. Sections 4, 5 and Appendix .1 provide several examples of a causing (or screening) variable $C$ in the context of Simpson’s paradox. The common cause principle was proposed to explain probabilistic correlations [34, 35]. It later found important applications in data science, where approximate relations similar to (15) are applied to effective data compression (Non-negative matrix factorization, Probabilistic Latent Dirichlet indexing, etc); see [40] for a review.
Note from (15) that $C$ gets rid of the conditional dependence on $B$ in $p(A_{1},A_{2}|B,C)$. Thus, a sensible way of looking at the association between $a_{1}$ and $a_{2}$ is to check the sign of
$p(a_{1}|a_{2},C)-p(a_{1}|\bar{a}_{2},C)~{}~{}~{}~{}{\rm for~{}each~{}value~{}of~{}}C.$ (17)
To support the usage of the common cause C for decision-making, we note that (15) has an important implication in the context of (1). (This implication generalizes the argument given in [35].) Assume that $p(a_{2},b,c)>0$ for all values $c$ of $C$. Note from (15) that there exists an event $c$ such that $p(a_{1}|a_{2},b)\leq p(a_{1}|a_{2},b,c)=p(a_{1}|a_{2},c)$, and an event $c^{\prime}$ such that $p(a_{1}|a_{2},b)\geq p(a_{1}|a_{2},b,c^{\prime})=p(a_{1}|a_{2},c^{\prime})$. Hence, if conditioning over $b$ facilitates (hinders) the association between $a_{1}$ and $a_{2}$, then conditioning over $c$ ($c^{\prime}$) is not worse in this facilitation (hindering)  [^4] To deduce the first relation assume that $p(a_{1}|a_{2},b)<p(a_{1}|a_{2},b,c)=p(a_{1}|a_{2},c)$ for all $c$, multiply both parts by $p(a_{2},b,c)>0$, sum over $c$ and get contradiction $p(a_{1},a_{2},b)<p(a_{1},a_{2},b)$. Likewise for the second relation..
@@
Finally, two remarks about the applicability of (1517). First, if $C$ is a common cause for both $A=(A_{1},A_{2})$ and $B$, the times of these variables naturally hold $t_{C}<{\rm min}[t_{A_{1}},t_{A_{2}},t_{B}]$. When $C$ screens $A$ from $B$, it holds $t_{B}<t_{C}<{\rm min}[t_{A_{1}},t_{A_{2}}]$. In certain applications of (17), it will suffice to have even a weaker condition $t_{C}<t_{A_{1}}$.
Second, we note that for applying (1, 2, 3) we do not need $p(A_{2})$, i.e. only $p(B|A_{2})$ is needed for connecting (1) with (2, 3). Indeed, $A_{2}$ does not necessarily need to be a random variable, but can simply be a label describing the situation. Now the same holds for (17): once (15) is written as
$p(A_{1},B|A_{2},C)=p(A_{1}|A_{2},C)p(B|C),$ (18)
we need only $p(C|A_{2})$ to pass from (18) to quantities involved in (1, 2, 3); i.e., $p(A_{2})$ is not needed.
3.2 A common cause (or screening variable) resolves Simpson’s paradox for binary variables
@@
Theorem 1: If $A_{1}$, $A_{2}$, $B$ and $C=\{c,\bar{c}\}$ are binary, and provided that (1) and (2, 3) are valid, all causes $C$ hold
@@
i.e. all $C$ holding (15) predict the same sign of association between $a_{1}$ and $a_{2}$ as (2, 3).
The main idea of proving (19) is inverting (15):
$p(a_{1}|a_{2},c)$
$=\frac{p(\bar{b}|\bar{c})p(a_{1}|a_{2},b)p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(a_{1}|a_{2},\bar{b})p(\bar{b}|a_{2})}{p(\bar{b}|\bar{c})p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(\bar{b}|a_{2})}$ (20)
@@
$p(c|a_{2})=\frac{p(\bar{b}|\bar{c})p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(\bar{b}|a_{2})}{p(b|c)+p(\bar{b}|\bar{c})-1}$
$=\frac{p(\bar{b}|\bar{c})+p(b|a_{2})-1}{p(b|c)+p(\bar{b}|\bar{c})-1},$ (22)
where unknown quantities $p(a_{1}|a_{2},c)$ and $p(c|a_{2})$ are represented via known ones (i.e. $p(A_{1},A_{2},B)$) and free parameters $p(B|C)$. Eqs. (21, 22) hold upon changing $a_{2}$ by $\bar{a}_{2}$ and are deduced in Appendix .3 via specific notations that should be useful when dealing with (15) for a non-binary $C$.
The rest of the proof is algebraic but non-trivial. It also works out and employs constraints (4, 31) on Simpson’s paradox itself. Expanding both sides of (1),
@@
and using there (2, 3) we subtract the sides of (1) from each other and find:
$p(a_{1}|a_{2},b)+p(\bar{b}|a_{2})[p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b)]<$
$p(a_{1}|\bar{a}_{2},b)+p(\bar{b}|\bar{a}_{2})[p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)].$ (25)
@@
Eqs. (23, 24) imply that for the validity of (13, 26) it is necessary to have $p(a_{1}|\bar{a}_{2},\bar{b})>p(a_{1}|a_{2},b)$, which together with (2, 3, 26) revert to (4). Now (1, 23, 24) read
$p(a_{1}|a_{2},b)+[1-p(b|a_{2})](p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))$ (27)
$<p(a_{1}|\bar{a}_{2},b)+[1-p(b|\bar{a}_{2})](p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$
$p(a_{1}|a_{2},\bar{b})-p(b|a_{2})(p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))$
$<p(a_{1}|\bar{a}_{2},\bar{b})-p(b|\bar{a}_{2})(p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$ (28)
where (27) and (28) are equivalent. Eqs. (27, 28, 4) imply
$[1-p(b|a_{2})](p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))<$
$[1-p(b|\bar{a}_{2})](p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$ (29)
$p(b|a_{2})(p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))>$
$p(b|\bar{a}_{2})(p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)).$ (30)
As checked directly, Eqs. (3.2, 30) lead to
@@
Now we return to (22) and assume there ${p(b|c)+p(\bar{b}|\bar{c})-1}<0$, which leads to ${p(\bar{b}|\bar{c})+p(b|a_{2})-1}<0$ from (22). Writing down from (22) the formula for $p(c|\bar{a}_{2})$ and making the same assumption we get ${p(\bar{b}|\bar{c})+p(b|\bar{a}_{2})-1}<0$. Now look at (21) and its analog obtained via $a_{2}\to\bar{a}_{2}$, and use there these two results together with (30, 31) and (4) to deduce the first inequality in (19) under assumption ${p(b|c)+p(\bar{b}|\bar{c})-1}<0$. It should be obvious that the second inequality in (19) holds under the same assumption since we nowhere used any specific feature of $c$ compared to $\bar{c}$.
@@
The rest is similar to the above: we proceed via (3.2, 31) and (4) and deduce (19) from (22), (3.2) and the analog of (3.2) obtained via $a_{1}\to\bar{a}_{2}$.
3.3 Non-binary causes
@@
Starting from a tertiary $C$, all three options of Simpson’s paradox become possible: there are common causes $C$ that support (1), those which support (2, 3), and eventually (random) cause variables $C=\{c_{1},c_{2},c_{3}\}$ for which $p(a_{1}|a_{2},c_{i})-p(a_{1}|\bar{a}_{2},c_{i})$ has different signs for different values of $i=1,2,3$. (Our numerical examples showing these possibilities are available upon request.) Hence, already for the tertiary cause one needs prior information on the common cause to decide on the solution of Simpson’s paradox. Alternatively, we can infer this unknown cause via e.g. one of the methods proposed recently [41, 42]. It is not excluded that such inference methods will provide further information on the solution of Simpson’s paradox. We hope to discuss this problem elsewhere.
4 Example: smoking and surviving
In sections 2.2.2 and 2.2.3 we discussed two examples studied in the literature and argued that they can be also interpreted via the common cause principle. In the present case, the standard approaches do not seem to apply, but the common cause can still be motivated. This example on survival of smokers versus nonsmokers is taken from Ref. [16]. Its technical details are discussed in Appendix .2.
Binary $A_{1}$ represents the survival in a group of women as determined by two surveys taken 20 years apart:
$A_{1}=\{a_{1},\bar{a}_{1}\}=\{{\rm died},{\rm alive}\},$
$A_{2}=\{a_{2},\bar{a}_{2}\}=\{{\rm smoker},{\rm nonsmoker}\},$ (33)
[…]
where $p(\bar{b})=0.1334$, and where $b$ and $\bar{b}$ denote age-groups. According to the data of [16], Simpson’s paradox reads
$p(a_{1}|a_{2})=0.2214<p(a_{1}|\bar{a}_{2})=0.2485,$ (35)
$p(a_{1}|a_{2},b)>p(a_{1}|\bar{a}_{2},b),$
$p(a_{1}|a_{2},\bar{b})>p(a_{1}|\bar{a}_{2},\bar{b}).$ (36)
Note that $B$ here influences $A_{1}$: the age of a person is a predictor of his/her survival. Causal influences from age to smoking can be neglected because the number of people that quit or started smoking is small [16]. We can assume that influences from smoking to age are absent. Then this example is intermediate between two situations considered in [7, 8, 9, 13]. Recall that when $B$ influenced $A_{2}$, these references advised to decide via the fine-grained option of the paradox, while for the case of the inverse influence (from $A_{2}$ to $B$) they recommend to employ the coarse-grained version; see Fig. 1.
Hence, we should expand on the above situation to achieve a workable model. We can assume that $A_{2}$ and $B$ are influenced by a common cause. Genetic factors influence an individual’s age and tendency to smoke. Originally proposed by Fisher [43], this hypothesis was later substantiated in several studies; see Refs. [44, 45] for reviews. Note that this refers to genetics of the smoking behavior itself, and not to health problems that can be caused by smoking plus genetic factors. Several sets of studies that contributed to genetic determinants of smoking behavior are as follows. (i) Children of smoking parents tend to smoke. (ii) Smoking behavior of adopted kids correlates stronger with that of their biological parents. (iii) Monozygotic (genetically identical) twins correlate in their smoking behavior much stronger than heterozygotic twins. Smoking behavior includes both the acquisition and maintenance of smoking. Monozygotic twins show correlations in both these aspects.
Hence as a preliminary hypothesis, we can suggest that genetic factors are the common cause of both smoking and age. If this common cause is binary, then Theorem 1 applies and we conclude—judging from the fine-grained data and consistently with other studies—that smoking is not beneficial for surviving.
5 Example: COVID-19, Italy versus China
Here the COVID-19 death rates are compared in Italy and China [31, 46]. According to the data, aggregated death rates in Italy are higher than in China, but in each age group, the death rates are higher in China. More precisely,
$A_{1}=\{a_{1},\bar{a}_{1}\}=\{{\rm died},{\rm alive}\},$
$A_{2}=\{a_{2},\bar{a}_{2}\}=\{{\rm China},{\rm Italy}\},$ (37)
[…]
where $p(a_{1})$ is the death rate out of COVID-19, $p(B)$ is found from the number of positively tested people in each age group, $p(\bar{b})=0.1012$, and where $p(\bar{b}|a_{2})=0.1017$ and $p(\bar{b}|\bar{a}_{2})=0.3141$. According to the data of [31], Simpson’s paradox reads
$p(a_{1}|a_{2})=0.0608<p(a_{1}|\bar{a}_{2})=0.0760,$ (39)
$p(a_{1}|a_{2},b)=0.0507>p(a_{1}|\bar{a}_{2},b)=0.04900,$
$p(a_{1}|a_{2},\bar{b})=0.150>p(a_{1}|\bar{a}_{2},\bar{b})=0.135.$ (40)
The authors of [31] proposed that this situation is described by TODAG $A_{2}\to B\to A_{1}\leftarrow A_{2}$; cf. (9). Then the conclusion from [9, 13] will be that the aggregated version of Simpson’s paradox works, i.e. Italy did worse than China. The authors of Ref. [31] reached the same conclusion.